The sum of the squares of two consecutive integers is 244. Find the two integers.
Answer by rfadrogane(214)

The sum of the squares of two consecutive integers is 244. Find the two integers.
Sol'n:think it two consecutive even integers, not a consecutive integers
let x-the first number
x+1-the second number
x^2+(x+2)^2 = 244
x^2+x^2+4x+4 = 244
2x^2+4x+4 = 244
2x^2+4x-240=0
by quadratic formula:
x=10 & -12
then,
if x = 10,x+2 = 12 ---answer
and
if x = -12, x+2 = -10 ----answer